# Stability analysis of a dark energy model in Rastall gravity

### v1

@@ Abstract
We study a cosmological model in Rastall’s theory of gravity in the framework of flat FLRW metric. We formulate the value of the Hubble parameter which contains two model parameters $\alpha$ and $j$. Employing the Markov Chain Monte Carlo (MCMC) sampling technique, we constrain the magnitude of the model parameters with their uncertainties. Moreover, we reconstruct the effective equation-of-state (EoS) parameter, which converges around the quintessence region. We perform a dynamical system analysis using the linearization technique to validate the results independently. Also, we discuss various physical properties of the model which shows the transition to acceleration and the violation of the Strong energy condition (SEC) in late times evolution. Finally, we conclude that our model mimics the behavior of a dark matter fluid during the past epoch and transitions into a quintessence dark energy model in later times.
@@ I Introduction
The General Theory of Relativity is an astounding accomplishment, and it is now widely considered to be one of the pillars of modern physics. The theory itself is couched in the language of differential geometry and was a pioneer for the use of modern mathematics in physical theories, leading the way for the gauge theories and string theories that have followed. It is no exaggeration to say that General Relativity set a new tone for what a physical theory can be and has truly revolutionized our understanding of the Universe. General Relativity has been established as the theory of gravitational interactions for over a century, being consistent with all experiments and being able to describe a huge set of observational results. Furthermore, there are still some shortcomings in theoretical and observational aspects of science, such as the accelerated universe and the existence of dark matter and energy, on account of which a modification of general relativity is inevitable.
There are possible modified and alternative theories of gravity that have been constructed in mainly two ways. In the first method, the fundamental assumptions of general relativity are still kept, but new additional terms are added to the Lagrangian density of general relativity, on account of which all field equations are modified with additional terms [1, 2, 3, 4, 5, 6, 7, 8]. In the second method, the fundamental assumptions of general relativity are changed. Rastall gravity belongs to the latter group. One of the fundamental assumptions of general relativity is the null covariant divergence of the energy-momentum tensor (EMT). However, Rastall followed the second method and argued that the vanishing divergence of the Einstein tensor does not necessarily imply the vanishing divergence of the energy-momentum tensor [9]. In such a scenario, energy conservation must be sacrificed. Relaxation of the condition of the vanishing of the divergence of the energy-momentum tensor does not upset the success of general relativity. Rastall gravity attempted to modify the theory of general relativity by assuming that the covariant derivative of the energy-momentum tensor varies directly as the derivative of the Ricci scalar [9]. It explicitly represents a non-minimal coupling between matter and geometry in curved spacetimes.
The one notable negative feature of Rastall gravity is the absence of Lagrangian. Recent progress in this direction indicates that a Lagrangian formulation for a Rastall-type theory can be given as an MG proposal where the Lagrangian depends on the Ricci scalar $R$ and the trace of the energy-momentum tensor $T$. Also, there are some questions regarding the status of Rastall gravity as an actual modification of general relativity and not only a redefinition of the energy-momentum tensor [10, 11]. At the geometric level, Rastall gravity is equivalent to Einstein gravity [11], unimodular theory as well as $f(R,T)$ theory; however, the point that the physical consequences are dramatically different has been amplified several times [12].
@@
Dynamical system analysis is an important tool in the study of cosmology, offering a method to qualitatively explore the behavior of solutions to cosmic models without drowning in complex calculations. Often, this method is employed to determine both the stability of the model and the existence of critical points [13, 14]. Through examining the asymptotic behavior of critical points of the dynamical system, one can explain the overarching dynamics of the universe across different cosmological epochs [15, 16, 17]. Researchers like Kadam et al. have presented an exploration of critical points representing different cosmic eras, including the matter-dominated, radiation-dominated, and dark energy eras in $f(T,B)$ gravity [18]. Stachowski and Szydłowski have tackled the dynamics of models by running cosmological constant terms with the help of a dynamical analysis approach [19]. Roy and Banerjee’s [20] extensive analysis of the stability of cosmological solutions for a spatially flat homogeneous and isotropic cosmological model within an extended Brans-Dicke theory through a phase space analysis serves as yet another testament to the power of dynamical system methods. Moreover, the versatility of dynamical systems extends to exploring interacting dark energy scenarios, offering a unified framework for understanding the cosmos at both background and perturbation levels [21, 22, 23, 24].
@@
The dynamical systems method is applied to probe critical points, and these points offer insights into the transitions between accelerating and decelerating epochs, characterizing phases dominated by dark energy, radiation, and matter [25]. Recent studies have promoted the discussion about the stability of various cosmological models using dynamical system analysis. For instance, Roy and Banerjee [26] scrutinized the stability criteria of chameleon models within the framework of spatially homogeneous and isotropic cosmology, employing the tools of autonomous system dynamical analysis. Similarly, Das et al. [27] explored the stability of the warm quintessential dark energy model. The collection of recent studies in modified gravity theories underscores the pivotal role of dynamical system analysis in advancing our understanding of the cosmos [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53].
@@
Over the last decades, Rastall gravity has gained the attention of scientists, and as a result, a vast number of papers have been published [11]. Furthermore, it incorporates a parameter that weights its deviations from general relativity. Over the last decade, the search for observational constraints on Rastall’s parameter has received increased attention, as has the study of applications to cosmology, black holes, and other compact objects. The properties of the solutions of black holes in Rastall gravity have been published [59, 60, 61, 62, 63, 64, 65, 66]. Rastall gravity offers some interesting results; for instance, de Sitter black holes have been found without explicitly assuming a cosmological constant [67]. It has been shown that Rastall gravity does admit the accelerated expansion of the universe. Various aspects of this theory, including theoretical and observational ones, have been reported in recent literature [68, 56, 25, 69].
@@
In this paper, we assemble our work in the following manner: In Sec. II, formulation for the Hubble parameter in terms of jerk parameter takes place. Afterward, in Sec. III, we obtained the Einstein field equations (EFEs) for flat FLRW space-time metric in Rastall gravity theory. Observational analysis has been done in Sec. IV, where three observational datasets are used to obtain the best-fitted values of model parameters. Sec. V comes along with the investigation of the physical and dynamical behavior of the universe. In Sec. VI, we accomplish a meticulous dynamical system analysis to discover the stability of the system depending on the various model parameters. In Sec. VII, we conclude our findings. Finally, in the Appendix. A, we consider the autonomous equation in 1D and determine the stability of the system around the fixed point.
II Formulation for the Hubble parameter in terms of jerk parameter
@@
In physics, jerk known as jolt is the rate of change of an object’s acceleration over time. It is a vector quantity. Jerk is denoted by the symbol $j$ and expressed in $m/s^{3}$ (SI units) or $g_{0}/s$ (standard gravities per second). As a vector, jerk $j$ can be expressed as the first-time derivative of acceleration, second-time derivative of velocity, and third-time derivative of position:
@@
where $q$, $v$, $r$, and $t$ represent acceleration, velocity, position and time respectively.
Third-order differential equations of the form
@@
are sometimes called jerk equations. When converted to an equivalent system of three ordinary first-order non-linear differential equations, jerk equations are the minimal setting for solutions showing chaotic behavior. This condition generates mathematical interest in jerk systems. Systems involving fourth-order derivatives or higher are accordingly called hyper-jerk systems [70].
If we define the jerk parameter $j$ as the function of redshift $z$, the relation between the jerk parameter $j$ and the deceleration parameter $q$ is given by
@@
The constant jerk parameter as $j=1$ corresponds to a flat $\Lambda$CDM model. Without loss of generality we consider the jerk parameter $j$ as an arbitrary constant [71, 72] and solving the differential equation (3), we get
@@
Rewriting Eq. (4) and using $\beta=1$ we get,
@@
By integrating both sides of the Eq. (7), we obtain:
@@
III Formulation of Einstein Field Equations in Rastall gravity
Rastall introduced a fact that the evolution of the universe does not perturb the rate of energy-momentum transference between matter fields and geometry. With this regard, Rastall supposed a coupling coefficient between the divergence of EMT and the derivative of the Ricci scalar, which is a constant and known as the Rastall parameter. This theory also describes the dark energy section to explain the accelerated universe. The generalized field equations for Rastall gravity theory are as follows
@@ III Formulation of Einstein Field Equations in Rastall gravity
where $\lambda$ is the Rastall parameter and $\kappa$ is the gravitational coupling constant. For flat, isotropic, and homogeneous universes the FLRW metric is written as
@@
where $a(t)$ stands for the scale factor of the universe. For the specific value of $\kappa$ and $\lambda$, we can come back to the theory of general relativity. Now, since Hubble parameter $H=\dot{a}/a$, therefore using Eqs. (10), (11), we get the Einstein field equations for Rastall gravity theory in terms of $H$ as
@@
where $\rho$ is the energy density, $p$ is the isotropic pressure and $\dot{H}$ indicates the derivative of $H$ with respect to cosmic time $t$.
@@
and the Equation-of-State parameter is given by $\omega(z)=p(z)/\rho(z)$, whose variations with respect to redshift $z$ can be seen in Fig. 4(c).
IV Observational analysis
Observational analysis is one of the advantageous parts of cosmology where we find the best-fitted results for the model parameters. In this analysis, we use the standard datasets to compute the optimal solution of the model parameters. Here, in this paper, we are using the Hubble dataset $H(z)$ (with $77$ points), $Pantheon$ dataset (with $1048$ points), and their joint dataset ($H(z)+Pantheon$), to calculate the best-fit value of $\alpha$ and $j$. For this purpose, we use the Markov Chain Monte Carlo (MCMC) method and the coding to plot the contour has been done in Python with the EMCEE library.
IV.1 $H(z)$ Dataset
$H(z)$ data acquired from cosmic chronometers which play an important role in studying the dark division of the universe [73]. In literature, various models have been discussed with different numbers of $H(z)$ data points [74, 75, 76, 77, 78]. In our model, we use $H(z)$ dataset along with $77$ data points [79]. We use the $\chi^{2}$ test to obtain the best-fit value of the model parameters and the formula for $\chi^{2}$ is written as
@@ IV.1 $H(z)$ Dataset
where $H(j,\alpha,z_{i})$ and $H_{obs}$ depict the theoretical and observed values of the Hubble parameter respectively, and $\sigma_{(z_{i})}$ shows the standard deviation for every $H(z_{i})$.
It has been observed that the analysis of the function $\chi_{Hb}^{2}$ using the Hubble dataset is not very authentic for any cosmological models. Therefore, we consider more reliable observation using the $Pantheon$ and their joint datasets.
Table 1: Best fit values of model parameters.
@@
IV.2 Pantheon Dataset
From the beginning of $21^{st}$ century, Type Ia supernovae are very useful for discussing the cosmological parameters as this is the first candidate to evidence the accelerated expansion of the Universe. Supernovae are stars and they they burst, they release a huge amount of energy and enlarge their outer shell. Thus, this part is related to examining the time evolution of their brightness and their spectrum. In our model, we use the $Pantheon$ datasets (1048 data points) and constrain the value of model parameters [80]. Pantheon dataset is the collection of data that have been compiled by different surveys e.g. the CfA1-CfA4 surveys, the Supernovae Legacy Survey (SNLS), the Pan-STARRS1 (PS1), the Carnegie Supernova Project (CSP), the Sloan Digital Sky Survey (SDSS), various Hubble Space Telescope (HST) etc. [81, 82, 83, 84, 85]. The range of redshift is between $0.01$ to $2.26$ in these surveys. In Type Ia supernovae analysis, we deal with redshift as well as luminosity distance which are closely related to standard candles in the Universe. The luminosity distance

### v2

@@ Abstract
We study a cosmological model in Rastall’s theory of gravity in the framework of the flat FLRW metric. We formulate the value of the Hubble parameter, which contains two model parameters, $\alpha$ and $j$. Employing the Markov Chain Monte Carlo (MCMC) sampling technique, we determine the values of these model parameters along with their uncertainties. Moreover, we derive the equation of state (EoS) parameter, which converges around the quintessence region. We perform a dynamical system analysis using the linearization technique to validate the results independently. Also, we discuss various physical attributes of the model, highlighting the transition to acceleration and the violation of the strong energy condition (SEC) in the late stages of evolution. In conclusion, our model mimics the behavior of a dark matter fluid during the past epoch and transitions into a quintessence dark energy model in the future epoch.
@@ I Introduction
The General Theory of Relativity stands as a monumental achievement and is universally acknowledged as a cornerstone of contemporary physics. This theory, articulated in the framework of differential geometry, pioneered the integration of advanced mathematics into physical theories, setting the stage for subsequent gauge and string theories [1, 2]. General relativity has redefined the expectations of a physical theory and has profoundly transformed our comprehension of the cosmos. For over a hundred years, it has stood as the absolute theory governing gravitational interactions, aligning with all experimental data and accounting for a vast array of observational findings [3, 4]. However, certain theoretical and observational challenges persist, such as the cosmic acceleration and the enigmatic presence of dark matter and energy, necessitating the consideration of potential modifications of general relativity [5, 6].
Various modifications and alternatives to the theory of gravity have been proposed, typically arising from two distinct methodologies. The first approach retains the fundamental postulates of general relativity while augmenting the Lagrangian density with additional terms, resulting in modified field equations [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. The second approach involves altering the fundamental assumptions of general relativity. Rastall gravity is an example of this latter category. Unlike the conventional perspective, which assumes a null covariant divergence of the energy-momentum tensor (EMT), Rastall proposed a different perspective, arguing that the zero divergence of the Einstein tensor does not necessarily lead to the zero divergence of the EMT [20]. This implies a departure from the principle of energy conservation. However, this relaxation of the zero divergence condition does not undermine the successes of general relativity. Rastall gravity sought to revise the theory of general relativity by positing that the covariant derivative of the EMT is proportional to the derivative of the Ricci scalar [20], thereby explicitly introducing a non-minimal coupling between matter and geometry in curved spacetimes.
A significant drawback of Rastall gravity lies in its absence of a Lagrangian formulation. However, some studies in the literature suggest a potential solution by formulating a Lagrangian-based model akin to Rastall theory as an MG model where the Lagrangian is dependent on the Ricci scalar $R$ and the trace of the energy-momentum tensor $T$. There are ongoing debates about whether Rastall gravity is a genuine modification of general relativity or merely a redefinition of the energy-momentum tensor [21, 22]. Geometrically, Rastall gravity aligns with Einstein’s gravity, unimodular theory, and $f(R,T)$ theory, but it’s been repeatedly emphasized that the physical implications are significantly different [23].
@@
Dynamical system analysis serves as a crucial tool in cosmological studies, offering a qualitative method to explore the behavior of solutions to cosmic models without drowning in complex calculations. Often, this method is employed to determine both the stability of the model and the existence of critical points [24, 25]. By examining the asymptotic behavior of critical points of the dynamical system, one can explain the overarching dynamics of the universe across different cosmological epochs [26, 27, 28]. Researchers like Kadam et al. have presented an exploration of critical points representing different cosmic eras, including the matter-dominated, radiation-dominated, and dark energy eras in $f(T,B)$ gravity [29]. Stachowski and Szydłowski have tackled the dynamics of models by running cosmological constant terms with the help of a dynamical analysis approach [30]. Roy and Banerjee’s [31] extensive analysis of the stability of cosmological solutions for a spatially flat homogeneous and isotropic cosmological model within an extended Brans-Dicke theory through a phase space analysis serves as yet another testament to the power of dynamical system methods. Moreover, the versatility of dynamical systems extends to exploring interacting dark energy scenarios, offering a unified framework for understanding the cosmos at both background and perturbation levels [32, 33, 34, 35].
@@
The dynamical systems method is applied to probe critical points, and these points offer insights into the transitions between accelerating and decelerating epochs, characterizing phases dominated by dark energy, radiation, and matter [36]. Recent studies have promoted the discussion about the stability of various cosmological models using dynamical system analysis. For instance, Roy and Banerjee [37] scrutinized the stability criteria of chameleon models within the context of spatially homogeneous and isotropic cosmology, employing the tools of autonomous system dynamical analysis. Similarly, Das et al. [38] explored the stability of the warm quintessential dark energy model. The collection of recent studies in modified gravity theories underscores the pivotal role of dynamical system analysis in advancing our understanding of the cosmos [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64].
@@
Over the past decades, Rastall gravity has attracted attention from researchers, resulting in a plethora of scholarly articles. Moreover, it includes a parameter that measures its deviations from general relativity. In the past ten years, there has been a notable surge in interest surrounding the exploration of observational constraints on Rastall’s parameter, along with the study of applications across various cosmological contexts, including black holes and other compact celestial bodies. The characteristics of the black hole solutions in Rastall gravity can be found in the literature [70, 71, 72, 73, 74, 75, 76, 77]. Rastall gravity offers some interesting results; for instance, de Sitter black holes have been found without explicitly assuming a cosmological constant [78]. It has been demonstrated that Rastall gravity accommodates the accelerated expansion of the universe. Several recently published literary works focus on different aspects of this theory [79, 67, 36, 80].
@@
In this paper, we assemble our work in the following manner: Sec. II comprises the brief background of Rastall gravity theory. In Sec. III, the formulation for the Hubble parameter in terms of the jerk parameter takes place. Afterward, in Sec. IV, we obtain the Einstein field equations (EFEs) for the flat FLRW space-time metric in Rastall gravity theory. Observational analysis has been done in Sec. V, where three observational datasets are used to obtain the best-fitted values of model parameters. Sec. VI comes along with the investigation of the physical and dynamical behavior of the universe. In Sec. VII, we accomplish a meticulous dynamical system analysis to discover the stability of the system depending on the various model parameters. In Sec. VIII, we conclude our findings. Finally, in the Appendix. A, we consider the autonomous equation in 1D and determine the stability of the system around the fixed point.
II Rastall Gravity
In this section, we aim to obtain a solution while taking into consideration the physical background of the gravity sector together with the matter source. We include two parts of the model: the Rastall field equations and the matter field equations, to start the computation in the Rastall gravity sector. In this modification of General Relativity (GR), the covariant conservation equation of the energy-momentum tensor expressed as $\nabla_{j}T^{ij}=0$ was changed to a more generalized version as [81]
$\nabla_{j}T^{ij}=u^{i}.$ (1)
To make the theory consistent with GR, the right-hand side of the Eq. (1) must be equal to zero when the scalar curvature or the background curvature vanishes. Thus, for our convenience, we can take the four-vector $u^{i}$ as,
$u^{i}=\lambda\nabla_{i}R,$ (2)
where $\lambda$ is a free parameter known as Rastall’s parameter. From Eqs. (1) and (2), we have the Rastall field equation as
$R_{ij}-\frac{1}{2}\left(1-2\beta\right)Rg_{ij}=\kappa T_{ij},$ (3)
where $\beta=\kappa\lambda$ and $\kappa$ is the gravitational coupling constant. The trace of the Eq. (3) yields
$R=\frac{\kappa}{4\beta-1}T,~{}\beta\neq 1/4.$ (4)
The field equation in the presence of a non-vanishing cosmological constant $\Lambda$ is
$E_{ij}+\Lambda g_{ij}+\beta g_{ij}R=\kappa T_{ij},$ (5)
where $E_{ij}$ is the standard Einstein tensor.
III Formulation for the Hubble parameter in terms of the jerk parameter
@@
In physics, the term jerk, also referred to as jolt, denotes the rate at which an object’s acceleration changes over time. It is a vector quantity and is symbolized by $j$. Its units are $m/s^{3}$ in the International System of Units (SI) or $g_{0}/s$ in terms of standard gravities per second. As a vector, jerk $j$ is mathematically defined as the first derivative of acceleration, the second derivative of velocity, and the third derivative of position with respect to time:
@@
where $q$, $v$, $r$, and $t$ represent acceleration, velocity, position, and time, respectively.
Differential equations of the third order,
@@
are colloquially termed jerk equations. When transformed into a system of three ordinary first-order non-linear differential equations, jerk equations represent the fundamental framework for modeling chaotic phenomena. This characteristic piques mathematical curiosity in jerk systems. Correspondingly, systems involving derivatives of the fourth order or higher are classified as hyper-jerk systems [82].
If we look at the jerk parameter, denoted by $j$, as a function of the redshift $z$, there exists a specific relationship between the jerk parameter $j$ and the deceleration parameter $q$, which is given by
@@
When the jerk parameter is constant and equals $1$, it corresponds to a flat $\Lambda$CDM model. Without loss of generality, we consider the jerk parameter $j$ as an arbitrary constant [83, 84] and solving the differential equation (8), we get
@@
Rewriting Eq. (9) and using $\beta=1,$ we get
@@
By performing integration on both sides of the Eq. (12), we obtain
@@
IV Formulation of Einstein Field Equations in Rastall Gravity
Rastall introduced the fact that the evolution of the universe does not perturb the rate of energy-momentum transference between matter fields and geometry. In this regard, Rastall supposed a coupling coefficient between the divergence of EMT and the derivative of the Ricci scalar, which is a constant and known as the Rastall parameter. This theory also describes the dark energy section to explain the accelerated universe. From Eq. (5), the generalized field equations for Rastall gravity theory in the absence of the cosmological constant $\Lambda$ are given by
@@ IV Formulation of Einstein Field Equations in Rastall Gravity
where $\lambda$ is the Rastall parameter and $\kappa$ is the gravitational coupling constant. For flat isotropic Universe, the FLRW metric is written as
@@
where $a(t)$ stands for the scale factor of the universe. For the specific values of $\kappa$ and $\lambda$, we can come back to the theory of general relativity. Now, since Hubble parameter $H=\dot{a}/a$, therefore using Eqs. (15), (16), we get the Einstein field equations for Rastall gravity theory in terms of $H$ as
@@
where $\rho$ is the energy density, $p$ is the isotropic pressure, and $\dot{H}$ indicates the derivative of $H$ with respect to cosmic time $t$.
@@
and the EoS parameter is given by $\omega(z)=p(z)/\rho(z)$, whose variations with respect to redshift $z$ can be seen in Fig. 4(c).
V Observational Analysis
Observational analysis is one of the advantageous parts of cosmology, enabling us to determine the best-fitted results for the parameters of the model. In this analysis, we use the standard datasets to compute the optimal solution of the model parameters. Here, in this paper, we are using the Hubble dataset $H(z)$ (with $77$ points), the $Pantheon$ dataset (with $1048$ points), and their joint dataset ($H(z)+Pantheon$), to calculate the best-fit value of $\alpha$ and $j$. For this purpose, we use the Markov Chain Monte Carlo (MCMC) method, and the coding to plot the contour has been done in Python with the EMCEE library.
V.1 $H(z)$ Dataset
$H(z)$ data acquired from cosmic chronometers, which play an important role in studying the dark division of the universe [85]. In literature, various models have been discussed with different numbers of $H(z)$ data points [86, 87, 88, 89, 90]. The presented model uses $H(z)$ dataset having $77$ data points [91]. We use the $\chi^{2}$ test to obtain the best-fit value of the model parameters, and the formula for $\chi^{2}$ is written as
@@ V.1 $H(z)$ Dataset
where $H(j,\alpha,z_{i})$ and $H_{obs}$ depict the theoretical and observed values of the Hubble parameter, respectively, and $\sigma_{(z_{i})}$ shows the standard deviation for every $H(z_{i})$.
The analysis of the function $\chi_{Hub}^{2}$ using the Hubble dataset has been found to be not very reliable for any cosmological models. Hence, we opt for more reliable observation using the $Pantheon$ and their joint datasets.
Table 1: Model Parameters: Best-Fit Results
@@
V.2 Pantheon Dataset
From the beginning of the $21^{st}$ century, Type Ia supernovae have been very useful for discussing the cosmological parameters, as this is the first candidate to evidence the accelerated expansion of the universe. Supernovae are stars that, upon explosion, emit a vast amount of energy and expand their outer shell. This aspect is related to the study of the time evolution of their spectrum and brightness. In our model, we use the $Pantheon$ dataset ($1048$ data points) and constrain the value of model parameters [92]. The $Pantheon$ dataset is the collection of data that have been compiled by different surveys, e.g., the CfA1-CfA4 surveys, the Supernovae Legacy Survey (SNLS), the Pan-STARRS1 (PS1), the Carnegie Supernova Project (CSP), the Sloan Digital Sky Survey (SDSS), various Hubble Space Telescope (HST) etc. [93, 94, 95, 96, 97]. The range of redshift is between $0.01$ and $2.26$ in these surveys. In the analysis of Type Ia supernovae, we deal with both redshift and luminosity distance, which are closely related to standard candles in the universe. The luminosity distance is